A wire carries a current of 10A in a direction 30º to a uniform magnetic field of
0.30 T. Find the force on 5.0 m of the wire.
1
Expert's answer
2020-05-24T17:59:17-0400

Given data

Current "I=10 A"

Magnetic field is "B=0.30 T"

Length of wire is "L=5.0 m"

The magnetic force on the wire is

"F=BILsin 30^o=(0.30)(10)(5.0)sin30^o=7.50 N"